Americans ate an average of 25.7 pounds of confectionery products each last year and

spent an average of $61.50 per person doing so. If the standard deviation of consumption is 3.75 pounds and the standard deviation of the amount spent is %5.89, find the following:


a. The probability that the sample mean confectionary consumption for a random sample

of 40 American consumers was greater than 27 pounds.


b. The probability that for a random sample of 50, the sample mean for confectionary

spending exceeded $60.00


1
Expert's answer
2022-06-10T05:30:15-0400

Solution:

Let's denote given values:

"\mu1=25.7pounds;"

"\mu2=61.5" $;

"\sigma1=3.75pounds;"

"X1=27pounds" -sample value for a);

"X2=60" $ - sample values for b);

"\sigma2=5.89" $;

"N1=40" - sample size for a);

"N2=50" - sample size for b);

We need to find z-scores:

"z=\frac{X-\mu}{\sigma_e};" here "\sigma_e=\frac{\sigma}{\sqrt{N}}" - standard deviation error when N sample numbers;

So,

a) "z1=\frac{X1-\mu1}{\sigma_e1};" "\sigma_{e1}=\frac{\sigma1}{\sqrt{N1}}=\frac{3.75}{\sqrt{40}}=0.59293;"

"z1=\frac{27-25.7}{0.59293}=2.19;"

Now, we need to check z-table, which "P=0.9857" for "z1=2.19."

If we want to find greater then 27 pounds, so

"P1=1-P=1-0.9857=0.0143;" or "P1=1.43" %.


b) "z2=\frac{X2-\mu2}{\sigma_e2};" "\sigma_{e2}=\frac{\sigma2}{\sqrt{N2}}=\frac{5.89}{\sqrt{50}}=0.833;"

"z2=\frac{60-61.5}{0.833}=-1.8;"

"P=0.0359" for "z2=-1.8."

If we want to find greater than 60$, so

"P2=1-P=1-0.0358=0.9641;" or "P2=96.41" %.

Answer:

a) "P1=1.43" %.

b) "P2=96.41" %.